# GATES: Cost-aware Dynamic Workflow Scheduling via Graph Attention Networks and Evolution Strategy

### v1

@@ Abstract
Cost-aware Dynamic Workflow Scheduling (CADWS) is a key challenge in cloud computing, focusing on devising an effective scheduling policy to efficiently schedule dynamically arriving workflow tasks, represented as Directed Acyclic Graphs (DAG), to suitable virtual machines (VMs). Deep reinforcement learning (DRL) has been widely employed for automated scheduling policy design. However, the performance of DRL is heavily influenced by the design of the problem-tailored policy network and is highly sensitive to hyperparameters and the design of reward feedback. Considering the above-mentioned issues, this study proposes a novel DRL method combining Graph Attention Networks-based policy network and Evolution Strategy, referred to as GATES. The contributions of GATES are summarized as follows: (1) GATES can capture the impact of current task scheduling on subsequent tasks by learning the topological relationships between tasks in a DAG. (2) GATES can learn the importance of each VM to ready tasks, increasing the chance of selecting the optimal VM. (3) Utilizing Evolution Strategy’s robustness, exploratory nature, and tolerance for delayed rewards, GATES achieves stable policy learning in CADWS. Extensive experimental results demonstrate the superiority of the proposed GATES in CADWS, outperforming several state-of-the-art algorithms. Codes are available at: https://github.com/YaShen998/GATES
@@ 1 Introduction
The key challenges of this study are: (1) how to design an effective policy network to properly control the influence of every scheduling decision on future task execution in a dynamic environment with continuously arriving workflows of highly diverse patterns; (2) how can the designed policy network support a time-varying collection of VMs; and (3) how to reliably train this policy network to ensure robust performance across varying system settings. To address the aforementioned challenges, this paper proposes a novel DRL method that integrates Graph Attention Networks (GAT) and Evolution Strategy (ES), referred to as GATES111“GATES” is used coincidentally and has no relation to Ning et al. (2020) in terms of problem setting or technical content..
Specifically, the main contributions of GATES are as follows: (1) GATES employs GAT to efficiently and scalably process the local topological information of each ready task. Leveraging GAT’s attention mechanism, GATES accurately learns the global attention information between the ready task and other tasks, enabling a holistic view of current scheduling impacts on future tasks by aggregating all task embeddings. (2) Dynamic heterogeneous graphs with identical structures but varying scales are constructed to represent the relationships between the ready task and all available VMs at each time step. Regardless of dynamic changes in VM numbers, GAT can compute attention coefficients from these graphs to identify the most suitable VM at each time step. (3) In CADWS, the reward function typically includes VM rental costs and SLA violation penalties. However, since SLA penalties can only be accurately determined after workflow execution, it’s difficult to design a reward function that quantifies SLA violations at each time step. To overcome this challenge, ES is utilized to train the policy network. This approach eliminates the need for complex reward design and hyperparameter tuning, while also preventing performance issues caused by poorly designed rewards.
@@ 2 Related Work
Recently, some studies Sun and Yang (2023); Zhang et al. (2024) proposed graph-based policy networks to extract state features for effective policy learning by leveraging the graph topological properties inherent in problems, such as traveling salesman, maximal independent set, and job shop problems. Sun and Yang [2023] proposed a graph-based diffusion model to exploit the graph topological information in the traveling salesman problem and the maximal independent set problem. Similarly, Zhang et al. [2024] utilized GNNs to extract state features for effective policy learning by leveraging the graph topological properties of the job shop scheduling problems. While the above studies considered the graph topological information, they assume static structure and fixed scheduling resources. In contrast, CADWS accounts not only for the dynamic workflow arrivals and patterns but also for the variability in scheduling resources over time Shen et al. (2024). Specifically, the pool of active VMs can change across different system states and time steps.
@@ 3.1 Preliminary of CADWS
The CADWS problem involves a set of workflows $\mathcal{W}$ arriving dynamically over time and a dynamically changing collection of heterogeneous VMs $\mathcal{M}$ for task execution, meaning that $|\mathcal{M}|$ is not fixed. Each workflow $W_{i}\in\mathcal{W}$ is modeled as a DAG (see Appendix A). $(\mathcal{O}_{W_{i}},\mathcal{C}_{W_{i}})$, where the node set $\mathcal{O}_{W_{i}}=\{O_{i1},O_{i2},\dots,O_{in}\}$ represents tasks, and the edge set $\mathcal{C}_{W_{i}}=\{(O_{ij},O_{ik})\mid O_{ij},O_{ik}\in\mathcal{O}_{W_{i}}\}$ denotes precedence constraints. An edge $(O_{ij},O_{ik})$ implies that task $O_{ij}$ must complete before task $O_{ik}$ can begin execution. $O_{ij}$ becomes a ready task $O^{*}$ waiting for current scheduling when there are no predecessor tasks for it.
@@ 4 Methology
Figure 1: The overall framework of CADWS
@@ 4.1 MDP Formulation of CADWS
System State: Each system state $s_{t}\in S$ represents a snapshot of the current state of the entire CADWS problem at decision step $t$, including the following elements: (1) the current ready task $O^{*}$ at $t$; (2) details of all active workflows, encompassing pattern information on their constituent tasks and task dependencies (modeled as Directed Acyclic Graphs, DAG), and workflow arrival times; (3) the numbers and processing status of all available VMs at $t$. By designing an effective policy network that accurately captures the three aforementioned factors, the system state of CADWS at any time step can be precisely represented, addressing the first two challenges mentioned in Section 1. To achieve this, we define the system state of CADWS at any time step as a DHG (as shown in Figure 2) and propose the HGAT employing GAT with a two-stage embedding mechanism to extract useful state representations for ready tasks (details in Figure 3), enabling the learning of effective scheduling policies and making informed decisions.
@@
Rewards: The reward $r_{t}$ provides a scalar feedback signal based on the action $a_{t}$ taken at system state $s_{t}$. To achieve the objective of minimizing the total cost as formulated in Eq. 1, the reward is defined as $r_{t}=-(VM_{Fee}+SLA_{Penalty})$.
@@
4.2 Dynamic Heterogeneous Graph
@@
Additionally, we incorporate dynamically changing VMs into the DAG to construct a Dynamic Heterogeneous Graph (DHG), as shown in Figure 2. Since the number and types of VMs vary dynamically at different decision steps, the VM nodes and edges in the DHG are inherently dynamic. The DHG provides a global perspective to describe the influence of all tasks and currently available VMs on the scheduling of the ready task. Moreover, the DHG is constructed specifically for each ready task: whenever a task becomes ready and awaits scheduling, a DHG is generated to capture the essential system’s dynamics for executing the ready task.
Figure 2: The proposed Dynamic Heterogeneous Graph (DHG)
@@ 4.3 Proposed Policy Network
To support scheduling policy learning in GATES (the Agent), it is crucial to extract state embedding from the constructed DHG. Graph Neural Networks (GNNs) are well-suited for learning embeddings from graph data. Among them, GAT outperforms Graph Convolutional Networks (GCN) and Graph Isomorphism Networks (GIN) in learning node embeddings from heterogeneous graphs. This advantage arises from GAT’s use of attention mechanisms to assign dynamic weights to each neighboring node, enabling it to adapt to heterogeneous node features and diverse neighborhood relationships Veličković et al. (2017). In contrast, GCN and GIN rely on fixed weights or mean aggregation, limiting their ability to model feature differences in heterogeneous graphs Thekumparampil et al. (2018). Moreover, GAT does not require strict alignment of node features and naturally handles various relationships in heterogeneous graphs through its multi-head attention mechanism, offering greater flexibility and scalability Wang et al. (2019).
@@ 4.3.1 Heterogeneous Graph Attention Networks (HGAT)
We propose a Heterogeneous Graph Attention Network (HGAT) to extract the state embedding from the DHG, as shown in Figure 3. The embedding learning process of HGAT consists of two stages: (1) Use GAT to learn feature embedding for each task from the DAG (Stage 1 of Figure 3), in which the Ready-task embedding 1 refers to the feature embedding of the ready task. Meanwhile, the feature embeddings of all tasks are aggregated using mean pooling to obtain the All tasks embedding; (2) Another GAT is employed to learn the Ready-task embedding 2 from the VM-graph in the DHG, as well as the VMs embeddings, enabling GATES to assess the relative importance of each VM to the ready task. Finally, these learned embeddings are concatenated to form the state embedding in GATES that comprehensively captures critical dynamics in CADWS, including the dynamic arrival of workflows, the changing DAG structures, and the dynamic variation in the number and status of VMs.
Figure 3: The proposed Heterogeneous Graph Attention Network (HGAT)
@@ 4.3.2 The Decision-making Networks
Figure 4: The decision-making networks
@@ 4.4 Raw features of CADWS
  • Number of successors: The count of tasks dependent on $O^{*}$.
@@
  • Task deadline: the task deadline is calculated using the method in Wu et al. (2017). The deadline is assigned to each task subject to the task size. A large task will be assigned an more relaxed task deadline.
@@ 4.5 Training process via Evolution Strategy
2. Fitness Evaluation: The fitness of each sampled parameter $\theta_{i}$ is evaluated as $F(\theta_{i})$, which corresponds to the total cost achieved by the policy $\pi_{i}$ on a given CADWS problem instance. The fitness function is defined as $F(\theta_{i})=F(\hat{\theta}+\sigma\epsilon_{i})=-\text{TotalCost}(\pi_{i})$, where $\text{TotalCost}(\pi_{i})$ represents the overall cost of executing $\pi_{i}$ as shown in Eq. 1.
@@
1:  while the current number of generation <=<=< = $Gen$ do
2:     Randomly sample a CADWS training instance: $Pro$.
3:     for each individual ($i$=1,2,…) in $N$ do
@@
6:        Evaluate the fitness value of $F(\theta_{i})$ using Eq. (1) based on $Pro$
@@ 5 Experiments
We conduct experiments using a dedicated simulator designed for the CADWS problem Yang et al. (2022, 2025), leveraging Amazon EC2-based virtual machine (VM) configurations and four workflow patterns/DAGs: CyberShake, Montage, Inspiral, and SIPHT, as shown in Appendix C. To reflect real-world resource heterogeneity, we employ six VM types with varying computational capacities and costs. Workflows are categorized into three scales, including small, medium, and large, based on the number of tasks, representing different levels of computational complexity. SLA deadlines are defined using a coefficient $\gamma\in\{1.00,1.25,1.50,1.75,2.00,2.25\}$, guiding resource selection under varying time constraints. More details on training and testing can be found in Appendix  D.
@@ 5.1 Experiment configuration
SLA Penalties: SLA deadlines were defined using $\gamma$, which influenced resource selection. SLA penalties encouraged renting cheaper VMs to manage cost while adhering to deadlines. The penalty weight $\beta$ was set to $0.24/hour$ Shen et al. (2024).
Baseline Algorithms: ProLis Wu et al. (2017) and GRP-HEFT Faragardi et al. (2019) are manual heuristic methods. ES-RL Huang et al. (2022) is a DRL-based scheduling policy, developed based on OpenAI’s code [^2] https://github.com/openai/evolution-strategies-starter. SPN-CWS Shen et al. (2024) is also a DRL-based method using the self-attention mechanism incorporated within the Transformer Vaswani (2017).
@@
In Figure 3, both stage 1 and stage 2 employ two GAT layers. The first layer has two attention heads and the second layer has one. The output dimensions $H$ of all GATs are 16. In Figure 4, the Transformer layers consist of 2 encoder layers, each with 2 self-attention heads, an input and output dimension of 16, and a hidden layer dimension of 128. In Stage 2 of the Decision Network, the MLP consists of 4 layers with hidden dimensions of 128 and output dimension of 1, and uses ReLU as the activation function. For ES in Algorithm 1, the hyperparameter settings are based on Huang et al. (2022) and Shen et al. (2024). Specifically, we set the population size to 40, the maximum number of generations to 2000, the initial learning rate to 0.01, and the standard deviation of Gaussian noise to 0.05.
@@ 5.2 Results and analysis
The results of all algorithms are presented in Table 3. The notation $\langle 1.00,S\rangle$ indicates that $\gamma=1.00$ and the algorithms are tested on the small-scale CADWS instance. ProLis and GRP-HEFT, being deterministic heuristic methods, do not have associated standard deviations. The symbols (+),(),(+),(-),( + ) , ( - ) , or ()($\approx$)( ≈ ) denote that the results are significantly better, worse, or statistically equivalent, respectively, compared to the corresponding algorithm. These comparisons are based on the Wilcoxon test with a significance level of 0.05.
@@
$\langle 1.00,S\rangle$ 773.69 1685.01(-) 223.68(51.51)(+)(+) 158.28(14.45)(+)(+)(+) 143.95(18.94)(+)(+)(+)(+)
$\langle 1.00,M\rangle$ 1829.25 2867.64(-) 422.26(100.92)(+)(+) 308.93(75.88)(+)(+)(+) 225.41(37.21)(+)(+)(+)(+)
$\langle 1.00,L\rangle$ 3641.79 5873.20(-) 829.64(262.91)(+)(+) 524.83(163.15)(+)(+)(+) 324.93(65.65)(+)(+)(+)(+)
$\langle 1.25,S\rangle$ 923.15 1967.31(-) 320.75(83.69)(+)(+) 143.12(15.73)(+)(+)(+) 124.14(7.10)(+)(+)(+)(+)
$\langle 1.25,M\rangle$ 2068.78 3578.04(-) 565.53(213.00)(+)(+) 272.77(60.65)(+)(+)(+) 193.26(24.06)(+)(+)(+)(+)
$\langle 1.25,L\rangle$ 4213.29 6868.22(-) 989.79(458.04)(+)(+) 496.63(155.02)(+)(+)(+) 300.19(76.29)(+)(+)(+)(+)
$\langle 1.50,S\rangle$ 901.40 1963.39(-) 270.93(67.20)(+)(+) 144.13(18.97)(+)(+)(+) 121.56(7.11)(+)(+)(+)(+)
$\langle 1.50,M\rangle$ 2035.52 3567.97(-) 505.44(222.45)(+)(+) 276.09(60.63)(+)(+)(+) 196.90(33.40)(+)(+)(+)(+)
$\langle 1.50,L\rangle$ 4066.84 6863.62(-) 886.87(380.26)(+)(+) 511.69(161.85)(+)(+)(+) 316.23(99.12)(+)(+)(+)(+)
$\langle 1.75,S\rangle$ 804.33 1959.71(-) 241.09(51.43)(+)(+) 137.36(14.10)(+)(+)(+) 118.45(4.85)(+)(+)(+)(+)
$\langle 1.75,M\rangle$ 1977.10 3560.22(-) 491.97(229.48)(+)(+) 263.12(57.61)(+)(+)(+) 195.25(35.83)(+)(+)(+)(+)
$\langle 1.75,L\rangle$ 3898.42 6854.40(-) 828.06(359.74)(+)(+) 502.67(155.64)(+)(+)(+) 315.84(92.48)(+)(+)(+)(+)
$\langle 2.00,S\rangle$ 789.71 1957.25(-) 225.70(50.89)(+)(+) 129.75(12.72)(+)(+)(+) 113.48(5.30)(+)(+)(+)(+)
$\langle 2.00,M\rangle$ 1921.42 3554.07(-) 427.16(213.71)(+)(+) 251.83(52.01)(+)(+)(+) 185.55(30.66)(+)(+)(+)(+)
$\langle 2.00,L\rangle$ 4024.03 6847.49(-) 780.77(408.14)(+)(+) 497.65(158.52)(+)(+)(+) 322.01(110.05)(+)(+)(+)(+)
$\langle 2.25,S\rangle$ 710.18 1954.71(-) 201.54(50.56)(+)(+) 126.18(13.02)(+)(+)(+) 110.90(5.70)(+)(+)(+)(+)
$\langle 2.25,M\rangle$ 1924.27 3551.31(-) 374.96(190.21)(+)(+) 238.35(48.26)(+)(+)(+) 179.17(24.54)(+)(+)(+)(+)
$\langle 2.25,L\rangle$ 3957.08 6842.88(-) 678.94(396.79)(+)(+) 475.46(139.99)(+)(+)(+) 318.19(112.21)(+)(+)(+)(+)
@@
5.3 Trade-off between VM rental fees and SLA penalty

### v3

@@ Abstract
Cost-aware Dynamic Workflow Scheduling (CADWS) is a key challenge in cloud computing, focusing on devising an effective scheduling policy to efficiently schedule dynamically arriving workflow tasks, represented as Directed Acyclic Graphs (DAG), to suitable virtual machines (VMs). Deep reinforcement learning (DRL) has been widely employed for automated scheduling policy design. However, the performance of DRL is heavily influenced by the design of the problem-tailored policy network and is highly sensitive to hyperparameters and the design of reward feedback. Considering the above-mentioned issues, this study proposes a novel DRL method combining Graph Attention Networks-based policy network and Evolution Strategy, referred to as GATES. The contributions of GATES are summarized as follows: (1) GATES can capture the impact of current task scheduling on subsequent tasks by learning the topological relationships between tasks in a DAG. (2) GATES can assess the importance of each VM to the ready task, enabling it to adapt to dynamically changing VM resources. (3) Utilizing Evolution Strategy’s robustness, exploratory nature, and tolerance for delayed rewards, GATES achieves stable policy learning in CADWS. Extensive experimental results demonstrate the superiority of the proposed GATES in CADWS, outperforming several state-of-the-art algorithms. The source code is available at: https://github.com/YaShen998/GATES.
@@ 1 Introduction
The key challenges of this study are: (1) how to design an effective policy network to properly control the influence of every scheduling decision on future task execution in a dynamic environment with continuously arriving workflows of highly diverse patterns; (2) how can the designed policy network support a time-varying collection of VMs; and (3) how to reliably train this policy network to ensure robust performance across varying system settings. To address the aforementioned challenges, this paper proposes a novel DRL method that integrates Graph Attention Networks (GAT) and Evolution Strategy (ES), referred to as GATES.
Specifically, the main contributions of GATES are as follows: (1) GATES employs GAT to efficiently and scalably process the local topological information of each ready task. Leveraging GAT’s attention mechanism, GATES accurately learns the global attention information between the ready task and other tasks, enabling a holistic view of current scheduling impacts on future tasks by aggregating all task embeddings. (2) Dynamic heterogeneous graphs with analogous structures but varying scales are constructed to represent the relationships between the ready task and all available VMs at each time step. Regardless of dynamic changes in VM numbers, GAT can compute attention coefficients from these graphs to identify the most suitable VM at each time step. (3) In CADWS, the reward function typically includes VM rental fees and SLA violation penalties. However, since SLA penalties can only be accurately determined after workflow execution, it’s difficult to design a reward function that quantifies SLA violations at each time step. To overcome this challenge, ES is utilized to train the policy network. This approach eliminates the need for complex reward design and hyperparameter tuning, while also preventing performance issues caused by poorly designed rewards.
@@ 2 Related Work
Recently, some studies Sun and Yang (2023); Zhang et al. (2024) proposed graph-based policy networks to extract state features for effective policy learning by leveraging the graph topological properties inherent in problems, such as traveling salesman, maximal independent set, and job shop problems. Sun and Yang [2023] proposed a graph-based diffusion model to exploit the graph topological information in the traveling salesman problem and the maximal independent set problem. Similarly, Zhang et al. [2024] utilized Graph Neural Networks (GNNs) to extract state features for effective policy learning by leveraging the graph topological properties of the job shop scheduling problems. While the above studies considered the graph topological information, they assume static structure and fixed scheduling resources. In contrast, CADWS accounts not only for the dynamic workflow arrivals and patterns but also for the variability in scheduling resources over time Shen et al. (2024). Specifically, the pool of active VMs can change across different system states and time steps.
@@ 3.1 Preliminary of CADWS
The CADWS problem involves a set of workflows $\mathcal{W}$ arriving dynamically over time and a dynamically changing collection of heterogeneous VMs $\mathcal{M}$ for task execution, meaning that $|\mathcal{M}|$ is not fixed. Each workflow $W_{i}\in\mathcal{W}$ is modeled as a DAG $(\mathcal{O}_{W_{i}},\mathcal{C}_{W_{i}})$, where the node set $\mathcal{O}_{W_{i}}=\{O_{i1},O_{i2},\dots,O_{in}\}$ represents tasks, and the edge set $\mathcal{C}_{W_{i}}=\{(O_{ij},O_{ik})\mid O_{ij},O_{ik}\in\mathcal{O}_{W_{i}}\}$ denotes precedence constraints (see Appendix A). An edge $(O_{ij},O_{ik})$ implies that task $O_{ij}$ must complete before task $O_{ik}$ can begin execution. $O_{ij}$ becomes a ready task $O^{*}$ waiting for current scheduling when there are no predecessor tasks for it Zhu et al. (2016).
@@ 4 Methology
Figure 1: The overall framework of CADWS.
@@ 4.1 MDP Formulation of CADWS
System State: Each system state $s_{t}\in S$ represents a snapshot of the current state of the entire CADWS problem at decision step $t$, including the following elements: (1) the current ready task $O^{*}$ at $t$; (2) details of all active workflows, encompassing pattern information on their constituent tasks and task dependencies (modeled as DAG), and workflow arrival times; (3) the numbers and processing status of all available VMs at $t$. By designing an effective policy network that accurately captures the three aforementioned factors, the system state of CADWS at any time step can be precisely represented, addressing the first two challenges mentioned in Section 1. To achieve this, we define the system state of CADWS at any time step as a DHG (as shown in Figure 2) and propose the HGAT employing GAT with a two-stage embedding mechanism to extract useful state embeddings for ready tasks (details in Figure 3), enabling the learning of effective scheduling policies and making informed decisions.
@@
Rewards: The reward $r_{t}$ provides a scalar feedback signal based on the action $a_{t}$ taken at system state $s_{t}$. To achieve the objective of minimizing the total cost as formulated in Eq. (1), the reward is defined as $r_{t}=-(VM_{Fee}+SLA_{Penalty})$.
@@
4.2 Dynamic Heterogeneous Graph (DHG)
@@
Additionally, we incorporate dynamically changing VMs into the DAG to construct a DHG, as shown in Figure 2. Since the number and types of VMs vary dynamically at different decision steps, the VM nodes and edges in the DHG are inherently dynamic. The DHG provides a global perspective to describe the influence of all tasks and currently available VMs on the scheduling of the ready task. Moreover, the DHG is constructed specifically for each ready task: whenever a task becomes ready and awaits scheduling, a DHG is generated to capture the essential system’s dynamics for executing the ready task.
Figure 2: The proposed Dynamic Heterogeneous Graph (DHG).
@@ 4.3 Proposed Policy Network
To support scheduling policy learning in GATES (the Agent), it is crucial to extract state embedding from the constructed DHG. GNNs are well-suited for learning embeddings from graph data. Among them, GAT outperforms Graph Convolutional Networks (GCN) and Graph Isomorphism Networks (GIN) in learning node embeddings from heterogeneous graphs. This advantage arises from GAT’s use of attention mechanisms to assign dynamic weights to each neighboring node, enabling it to adapt to heterogeneous node features and diverse neighborhood relationships Veličković et al. (2017). In contrast, GCN and GIN rely on fixed weights or mean aggregation, limiting their ability to model feature differences in heterogeneous graphs Thekumparampil et al. (2018). Moreover, GAT does not require strict alignment of node features and naturally handles various relationships in heterogeneous graphs through its multi-head attention mechanism, offering greater flexibility and scalability Wang et al. (2019).
@@ 4.3.1 Heterogeneous Graph Attention Networks (HGAT)
We propose an HGAT to extract the state embedding from the DHG, as shown in Figure 3. The embedding learning process of HGAT consists of two stages: (1) Use GAT to learn feature embedding for each task from the DAG (Stage 1 of Figure 3), in which the Ready-task embedding 1 refers to the feature embedding of the ready task. Meanwhile, the feature embeddings of all tasks are aggregated using mean pooling to obtain the All tasks embedding; (2) Another GAT is employed to learn the Ready-task embedding 2 from the VM-graph in the DHG, as well as the VMs embeddings, enabling GATES to assess the relative importance of each VM to the ready task. Finally, these learned embeddings are concatenated to form the state embedding in GATES that comprehensively captures critical dynamics in CADWS, including the dynamic arrival of workflows, the changing DAG structures, and the dynamic variation in the number and status of VMs.
Figure 3: The proposed HGAT: (a) a DHG tailored to the ready task, (b) the embedding learning process, (c) feature concatenation, (d) output the learned state embedding.
@@ 4.3.2 The Decision-making Networks
Figure 4: The decision-making networks.
@@ 4.4 Raw Features of CADWS
  • Forthcoming successors: the count of tasks dependent on $O^{*}$.
@@
  • Task deadline: the task deadline is calculated using the method in Wu et al. (2017). The deadline is assigned to each task, subject to the task size. A large task will be assigned a more relaxed task deadline.
@@ 4.5 Training Process via Evolution Strategy
2. Fitness Evaluation: The fitness of each sampled parameter $\theta_{i}$ is evaluated as $F(\theta_{i})$, which corresponds to the total cost achieved by the policy $\pi_{i}$ on a given CADWS problem instance. The fitness function is defined as $F(\theta_{i})=F(\hat{\theta}+\sigma\epsilon_{i})=-\text{TotalCost}(\pi_{i})$, where $\text{TotalCost}(\pi_{i})$ represents the overall cost of executing $\pi_{i}$ as shown in Eq. (1).
@@
1:while the current number of generation <=<= GenGen do
2:  Randomly sample a CADWS training instance: ProPro.
3:  for each individual (ii=1,2,…) in NN do
@@
6:   Evaluate the fitness value of $F(\theta_{i})$ using Eq. (1) based on ProPro
@@ 5 Experiments
We conduct experiments using a dedicated simulator designed for the CADWS problem Yang et al. (2022, 2025), leveraging Amazon EC2-based virtual machine (VM) configurations and four workflow patterns/DAGs: CyberShake, Montage, Inspiral, and SIPHT Deelman et al. (2015), as shown in Appendix C. To reflect real-world resource heterogeneity, we employ six VM types with varying computational capacities and costs. Workflows are categorized into three scales, including small, medium, and large, based on the number of tasks, representing different levels of computational complexity. SLA deadlines are defined using a coefficient $\gamma\in\{1.00,1.25,1.50,1.75,2.00,2.25\}$, guiding resource selection under varying time constraints. More details on training and testing can be found in Appendix D.
@@ 5.1 Experiment Configuration
SLA Penalties: SLA deadlines were defined using $\gamma$, which influenced resource selection. Larger SLA penalties encouraged renting cheaper VMs to manage cost while adhering to deadlines. The penalty weight $\beta$ was set to $0.24/hour$ Shen et al. (2024).
Baseline Algorithms: ProLis Wu et al. (2017) and GRP-HEFT Faragardi et al. (2019) are heuristic methods. ES-RL Huang et al. (2022) is a DRL-based scheduling policy, developed based on OpenAI’s code [^2] https://github.com/openai/evolution-strategies-starter. SPN-CWS Shen et al. (2024) is also a DRL-based method using the self-attention mechanism incorporated within the Transformer Vaswani (2017).
@@
In Figure 3, both stage 1 and stage 2 employ two GAT layers. The first layer has two attention heads and the second layer has one. The output dimensions $H$ of all GATs are 16. In Figure 4, the Transformer layers consist of 2 encoder layers, each with 2 self-attention heads, an input and output dimension of 16, and a hidden layer dimension of 128. In Stage 2 of the decision-making network, the MLP consists of 4 layers with hidden dimensions of 128 and output dimension of 1, and uses ReLU as the activation function. For ES in Algorithm 1, the hyperparameter settings are based on Huang et al. (2022) and Shen et al. (2024). Specifically, we set the population size to 40, the maximum number of generations to 2000, the initial learning rate to 0.01, and the standard deviation of Gaussian noise to 0.05.
@@ 5.2 Results and Analysis
The results of all algorithms are presented in Table 3. The notation $\langle 1.00,S\rangle$ indicates that $\gamma=1.00$ and the algorithms are tested on the small-scale CADWS instance. ProLis and GRP-HEFT, being deterministic heuristic methods, do not have associated standard deviations. The symbols (+),(),(+),(-), or $(\approx)$ denote that the results are significantly better, worse, or statistically equivalent, respectively, compared to the corresponding algorithm. These comparisons are based on the Wilcoxon test with a significance level of 0.05.
@@
$\langle 1.00,S\rangle$ 773.69 1685.01(-) 223.68(51.51)$(+)$ 158.28(14.45)(+)$(+)$ 143.95(18.94)$(+)(+)$
$\langle 1.00,M\rangle$ 1829.25 2867.64(-) 422.26(100.92)$(+)$ 308.93(75.88)(+)$(+)$ 225.41(37.21)$(+)(+)$
$\langle 1.00,L\rangle$ 3641.79 5873.20(-) 829.64(262.91)$(+)$ 524.83(163.15)(+)$(+)$ 324.93(65.65)$(+)(+)$
$\langle 1.25,S\rangle$ 923.15 1967.31(-) 320.75(83.69)$(+)$ 143.12(15.73)(+)$(+)$ 124.14(7.10)$(+)(+)$
$\langle 1.25,M\rangle$ 2068.78 3578.04(-) 565.53(213.00)$(+)$ 272.77(60.65)(+)$(+)$ 193.26(24.06)$(+)(+)$
$\langle 1.25,L\rangle$ 4213.29 6868.22(-) 989.79(458.04)$(+)$ 496.63(155.02)(+)$(+)$ 300.19(76.29)$(+)(+)$
$\langle 1.50,S\rangle$ 901.40 1963.39(-) 270.93(67.20)$(+)$ 144.13(18.97)(+)$(+)$ 121.56(7.11)$(+)(+)$
$\langle 1.50,M\rangle$ 2035.52 3567.97(-) 505.44(222.45)$(+)$ 276.09(60.63)(+)$(+)$ 196.90(33.40)$(+)(+)$
$\langle 1.50,L\rangle$ 4066.84 6863.62(-) 886.87(380.26)$(+)$ 511.69(161.85)(+)$(+)$ 316.23(99.12)$(+)(+)$
$\langle 1.75,S\rangle$ 804.33 1959.71(-) 241.09(51.43)$(+)$ 137.36(14.10)(+)$(+)$ 118.45(4.85)$(+)(+)$
$\langle 1.75,M\rangle$ 1977.10 3560.22(-) 491.97(229.48)$(+)$ 263.12(57.61)(+)$(+)$ 195.25(35.83)$(+)(+)$
$\langle 1.75,L\rangle$ 3898.42 6854.40(-) 828.06(359.74)$(+)$ 502.67(155.64)(+)$(+)$ 315.84(92.48)$(+)(+)$
$\langle 2.00,S\rangle$ 789.71 1957.25(-) 225.70(50.89)$(+)$ 129.75(12.72)(+)$(+)$ 113.48(5.30)$(+)(+)$
$\langle 2.00,M\rangle$ 1921.42 3554.07(-) 427.16(213.71)$(+)$ 251.83(52.01)(+)$(+)$ 185.55(30.66)$(+)(+)$
$\langle 2.00,L\rangle$ 4024.03 6847.49(-) 780.77(408.14)$(+)$ 497.65(158.52)(+)$(+)$ 322.01(110.05)$(+)(+)$
$\langle 2.25,S\rangle$ 710.18 1954.71(-) 201.54(50.56)$(+)$ 126.18(13.02)(+)$(+)$ 110.90(5.70)$(+)(+)$
$\langle 2.25,M\rangle$ 1924.27 3551.31(-) 374.96(190.21)$(+)$ 238.35(48.26)(+)$(+)$ 179.17(24.54)$(+)(+)$
$\langle 2.25,L\rangle$ 3957.08 6842.88(-) 678.94(396.79)$(+)$ 475.46(139.99)(+)$(+)$ 318.19(112.21)$(+)(+)$
@@
5.3 Trade-off between VM Fees and SLA Penalties